# Comparative Study of Quantum Transpilers: Evaluating the Performance of qiskit-braket-provider, qBraid-SDK, and Pytket Extensions

### v1

@@ Abstract
In this study, we conducted a comprehensive evaluation of several quantum transpilers, including the qiskit-braket-provider, the qBraid-SDK, and the pytket extensions, focusing on critical metrics such as correctness, failure rate, and transpilation time. Our results demonstrate that the qiskit-braket-provider exhibits superior performance achieving a remarkably low failure rate of 0.2%. the qiskit-braket-provider utilizes a combination of one-to-one transpilation and gate decomposition for unsupported gates, enhancing transpiler capabilities and speed. The qBraid-SDK offers a more generalized approach, suitable for transpilation across multiple SDKs, albeit with slower performance compared to the qiskit-braket-provider. The pytket extensions, while fast, exhibit limitations in handling complex circuits due to their one-to-one transpilation scheme. We also provide recommendations for future development, advocating for the adoption of the method employed by the qiskit-braket-provider to enhance transpiler capabilities and speed. This study contributes to the growing body of knowledge in quantum transpiler benchmarking, fostering interoperability and guiding the development of quantum computing applications in a diverse hardware and software landscape.
@@ I Introduction
Quantum computing holds the promise to transform our computational power and revolutionize our problem-solving capabilities[1, 2, 3]. In a relentless race toward practical implementations, companies strive to construct large, cost-effective, and robust quantum computing platforms. To this end, technologies like superconductors[4], trapped ions[5], neutral atoms[6], photonics[7], etc. are harnessed, each presenting unique advantages and challenges. Amidst the hardware race, a parallel effort unfolds in software development. Abstraction of physics inherent to diverse hardware platforms and streamlining quantum software development necessitate the creation of programming languages and software development kits (SDKs). Various SDKs have emerged to facilitate quantum programming such as IBM’s Qiskit[8], Google’s Cirq[9], Amazon’s Braket[10], Rigetti’s Quil[11], Xanadu’s Pennylane[12], and Quantinuum’s Pytket[13], among others. As the field is still in its nascent stage, each company’s focal points, principles, and applications are different, leading to each SDK carrying its own philosophy, capabilities, and limitations. Recognizing the impracticality of rewriting code for each SDK and for each quantum computer. Quantum transpilers have emerged as a crucial bridge in the quantum computing ecosystem. These transpilers enable users to craft quantum circuits using their SDK of choice and seamlessly convert them into equivalent circuits compatible with different SDKs. This allows users to exploit the unique capabilities of each SDK and execute circuits on multiple quantum computers without rewriting them. This paper aims to contribute to the expanding body of knowledge in quantum computing by addressing challenges and gaps in quantum transpilers benchmarking. Our goal is to provide a comprehensive evaluation of various quantum transpilers, considering critical metrics such as accuracy and speed. Through this analysis, we aspire to extract valuable insights that will not only enhance our understanding of transpiler performance but also guide improvements, thereby advancing the field. Additionally, our work aims to foster interoperability and guide the development of quantum computing applications in an era defined by diversity in both hardware and software.
This paper is structured as follows: In the subsequent section, we will elaborate on the methodology employed to benchmark various quantum transpilers. Section III will then showcase our results, and in section IV, we will provide a concise conclusion along with our recommendations.
@@ Acknowledgments
The current work was carried out during a remote internship at qBraid. We would like to express our gratitude to the qBraid team for their financial and technical support, guidance, and valuable insights throughout the duration of this research. Their expertise and resources significantly contributed to the success of this study.

### v2

@@ Abstract
In this study, we present a comprehensive evaluation of popular SDK-to-SDK quantum transpilers (that is transpilers that takes a quantum circuit from an initial SDK and output a quantum circuit in another SDK), focusing on critical metrics such as correctness, failure rate, and transpilation time. To ensure unbiased evaluation and accommodate diverse quantum computing scenarios, we developed two dedicated tools: RandomQC, for generating random quantum circuits across various types (pure random, VQE-like, and SDK-specific circuits), and Benchmarq, to streamline the benchmarking process. Using these tools, we benchmarked prominent quantum transpilers as of February 2024. Our results highlight the superior performance of the qiskit-braket-provider, a specialized transpiler from Qiskit to Braket, achieving a remarkably low failure rate of 0.2%. The qBraid-SDK, offering generalized transpilation across multiple SDKs, demonstrated robust but slower performance. The pytket extensions, while fast, faced limitations with complex circuits due to their one-to-one transpilation approach. In particular, the exceptional performance of the qiskit-bracket-provider stems not only from its specialization but also from its architecture, which combines one-to-one transpilation with gate decomposition for unsupported gates, enhancing both speed and capability. This study aims to provide practical guidelines to users of SDK-to-SDK quantum transpilers and guidance to developers for improving the design and development of future tools.
@@ I Introduction
Quantum computing holds the promise to transform our computational power and revolutionize our problem-solving capabilities[1, 2, 3]. In a relentless race toward practical implementations, companies strive to construct large, cost-effective, and robust quantum computing platforms. To this end, technologies such as superconductors[4], trapped ions[5], neutral atoms[6], photonics[7], etc. are harnessed, each presenting unique advantages and challenges. Amidst the hardware race, a parallel effort unfolds in software development. Abstraction of physics inherent to diverse hardware platforms and streamlining quantum software development necessitate the creation of programming languages and software development kits (SDKs). Various SDKs have emerged to facilitate quantum programming such as IBM’s Qiskit[8], Google’s Cirq[9], Amazon’s Braket[10], Rigetti’s Quil[11], Xanadu’s Pennylane[12], and Quantinuum’s Pytket[13], among others. As the field is still in its nascent stage, each company’s focal points, principles, and applications are different, leading to each SDK carrying its own philosophy, capabilities, and limitations. Recognizing the impracticality of rewriting code for each SDK and for each quantum computer. SDK-to-SDK quantum transpilers have emerged as a crucial bridge in the quantum computing ecosystem. These transpilers enable users to craft quantum circuits using their SDK of choice and seamlessly convert them into equivalent circuits compatible with different SDKs. This allows users to exploit the unique capabilities of each SDK and execute circuits on multiple quantum computers without rewriting them.
This study aims to provide practical guidelines to users of SDK-to-SDK quantum transpilers and guidance to developers for improving the design and development of future tools. At the time of writing this article (Q1 2024), we are not aware of any similar study that comprehensively evaluates SDK-to-SDK quantum transpilers.
This paper is structured as follows. In the subsequent section, we will elaborate on the methodology employed to benchmark various quantum transpilers. Section III will then showcase our results, and in Section IV, we will provide a concise conclusion along with our recommendations.
@@ Acknowledgments
The current work was carried out during a remote internship at qBraid. We would like to express our gratitude to the qBraid team for their financial and technical support, guidance, and valuable insights throughout the duration of this research. Their expertise and resources significantly contributed to the success of this study. This document has been produced with the financial assistance of the European Union (Grant no. DCI-PANAF/2020/420-028), through the African Research Initiative for Scientific Excellence (ARISE), pilot programme. ARISE is implemented by the African Academy of Sciences with support from the European Commission and the African Union Commission. The contents of this document are the sole responsibility of the author(s) and can under no circumstances be regarded as reflecting the position of the European Union, the African Academy of Sciences, and the African Union Commission. We are grateful to the Algerian Ministry of Higher Education and Scientific Research and DGRST for the financial support.